# The gravitational-wave emission from the explosion of a 15 solar mass star with rotation and magnetic fields

### v1

@@ 4 Gravitational Wave Properties
Figure 5: The time series of the gravitational-wave signal from matter for all models as measured for an observer at the equator ($\theta=\phi=90$). The left column shows the plus polarisation, and the right column shows the cross polarisation. The rapidly rotating models have higher gravitational-wave amplitudes. Some of the models have started to develop low-frequency tails before the end of the simulation time.
The time series of the gravitational waves from matter as measured by an observer in the equatorial plane ($\theta=\phi=90^{\circ}$), are shown in Figure 5. The gravitational-wave signals are upgraded from 2D to 3D at post-bounce times of 5 ms for SFHx_rr, 3 ms for SFHo_rr, 31 ms for SFHx_nr and 13 ms for SFHo_nr. As the core-bounce time was simulated in 2D, the non-rotating models are showing strong ringdown after prompt convection at the start, and slowly evolve to lower gravitational-wave amplitudes as the initial axisymmetry is lost in favour of less coherent motions in 3D. This has the largest impact on model SFHo_nr, which has a high amplitude of 10 cm at the start in the plus polarisation. This issue does not have an impact on the rotating models where there is a (physically) preferred axis anyway.
The non-rotating models reach gravitational-wave amplitudes of $\sim 10$ cm. This is similar to typical amplitudes from our previous neutrino-driven explosion models, for example, model s18 from Powell & Müller (2019), and model y20 from Powell & Müller (2020). This shows that the addition of weak magnetic fields does not have a big impact on gravitational-wave amplitudes. The rotating models reached gravitational-wave amplitudes of $\sim 40$ cm. This amplitude is similar to our previous rapidly rotating model m39 (Powell & Müller, 2020), which had no magnetic fields, but significantly smaller than our previous models with strong magnetic fields and rapid rotation, models m39_B10 and m39_B12 from Powell et al. (2023). This is consistent with the much lower explosion energy in the current models compared to $\sim 3\times 10^{51}\,\mathrm{erg}$ in m39_B10 and m39_B12, and suggests that a stronger magnetic field strength is required for powerful gravitational-wave signals from magnetorotational explosions.
@@
Figure 6: Spectrograms of $h_{+}+h_{\times}$ for all our gravitational-wave signals. The top left panel shows model SFHo_rr, the top right panel shows model SFHo_nr, the bottom left panel shows model SFHx_rr, and the bottom right panel shows model SFHx_nr. The non-rotating models show some low frequency gravitational-wave emission from the SASI before shock revival time. All models reach high frequencies of above 1000 Hz before the end of the simulation. Model SFHx_rr reaches higher frequencies due to its smaller PNS radius.
Spectrograms of the gravitational-wave emission are given in Figure 6. At early times, the non-rotating models have a high-power hotspot at low frequency caused by the ringdown after the transition from 2D to 3D. In the non-rotating models, some low-frequency emission due to the SASI is visible before 0.2 s when shock revival stops the SASI activity. The non-rotating models both have a similar high-frequency f/g-mode emission. They both reach a maximum frequency of $\sim 1500$ Hz by the end of the simulation. The gravitational-wave frequency is higher than in our previous simulations because of the pseudo-Newtonian gravity used in our simulation. Previous studies have shown that the gravitational-wave frequency is about $\sim 20$% higher than when full general relativity is used (Müller et al., 2013), due to missing relativistic correction terms in the Brunt-Väisälä frequency in the pseudo-Newtonian approximation. The rapidly rotating models have no visible lower frequency modes. Their high-frequency emission has a different shape and maximum frequency for the two different EoS, with a slight S-shape for model SFHx_rr. This is due to the difference in the PNS radius between the two models, as shown previously in Figure 1, where the smaller final PNS radius for model SFHx_rr has resulted in a higher maximum frequency of the f/g-mode mode. The dominant emission band is distinctly broader for the models with rapid rotation, with a width of $\sim 500$ Hz instead of $\sim 250$ Hz, reminiscent of the cases with rapid rotation and strong fields in the 2D study of (Jardine et al., 2022). Such a broadening of the dominant emission band may be an indicator for rapid rotation in the event of a gravitational-wave detection from a Galactic CCSN.
@@ 5 The Neutrino Memory
Figure 7: The gravitational-wave amplitude time series from both matter and anisotropic neutrino emission. The left column shows the plus polarisation, and the right column shows the cross polarisation, both calculated at $\theta=\phi=90^{\circ}$ (equatorial observer). All models show a significant increase in amplitude at low frequencies (slow variations) compared to the matter signal due to the neutrino emission anisotropy.
In addition to the low-frequency component to the gravitational-wave emission from matter, there generally is a high-amplitude low-frequency component to the gravitational-wave emission from the asymmetric emission of neutrinos (neutrino memory, Epstein, 1978). The neutrino memory will usually dominate over the low-frequency matter signal (Marek et al., 2009b; Müller et al., 2013). This component of the gravitational-wave signal is often ignored in CCSN simulations that focus on gravitational waves, as it is outside of the frequency band for the current gravitational-wave detectors. However, understanding the gravitational-wave signals at frequencies below 10 Hz is important for the science case for the next generation of gravitational-wave detectors. In this section, we show how the gravitational-wave emission changes when the signal from the asymmetric emission of neutrinos is also included.
To calculate the gravitational-wave signal from anisotropic emission of neutrinos, we use Equation (24) from Mueller & Janka (1997), which gives the transverse-traceless perturbations $h_{ij}^{\mathrm{TT}}$ of the metric from neutrino memory,
@@
The combined gravitational-wave emission from both matter and neutrinos is shown in Figure 7. The low-frequency gravitational-wave signal from asymmetric neutrino emission has significantly higher amplitude than the matter signal alone. All models show a significant increase in gravitational-wave amplitude when the neutrino memory is included. Model SFHo_rr reaches amplitudes of 200 cm in the plus polarization, model SFHo_nr reaches 140 cm, model SFHx_rr reaches 100 cm, and model SFHo_nr reaches -150 cm. Due to the short simulation time of these models, we can only predict the low-frequency emission down to about 2 Hz, as the minimum measurable frequency is determined by the simulation duration. The fact that the models quickly reach high amplitudes at low frequencies reflects that the models develop large-scale asymmetry in the explosion ejecta soon after the shock is revived. There is slightly less anisotropy in model SFHx_rr than in the others. The gravitational-wave signal is still growing significantly in time for the non-rotating models at the end of the simulation time. Longer duration simulations by other groups have shown that the neutrino memory can grow to over a 1000 cm at later times (Vartanyan et al., 2023). This shows the need to carry out long duration simulations to fully capture the both the minimum frequency and maximum amplitude of low-frequency gravitational-wave emission.
We also show the amplitude spectral densities for all models, with the matter components only, and both the matter and neutrino components in Figure 8. Including the neutrino memory results in a much larger amplitude at low frequencies, in the frequency detection bands of Einstein Telescope and proposed moon-based gravitational-wave detectors.
Figure 8: The amplitude spectral density (ASD) of all models, as measured at a distance of 100 kpc, and the sensitivity curves of the LIGO, Einstein Telescope (ET), Cosmic Explorer and moon-based (GLOC) gravitational-wave detectors. The left panel shows the matter component only. The right panel shows the spectrum from the combined matter and neutrino signal. The gravitational wave spectrum from matter peaks at frequencies of above 1000 Hz, and the gravitational wave spectrum from asymmetric neutrino emission peaks below 10 Hz.
@@ 6 Gravitational Wave Directional Dependence and detectability
Figure 9: The gravitational-wave time series of all four models as measured at six different observer angles. (top left) The gravitational waves for model SFHo_rr. (top right) The gravitational waves for model SFHo_nr. (bottom left) The gravitational waves for model SFHx_rr. (bottom right) The gravitational waves for model SFHx_nr. The amplitude of the low-frequency gravitational-wave emission can vary significantly at different source angles.
In Figure 9, we show the gravitational-wave emission calculated at six different representative source angles $(\theta,\phi)\in\{(0^{\circ},0^{\circ}),(180^{\circ},0^{\circ}),(90^{\circ},0^{\circ}),(90^{\circ},90^{\circ}),(90^{\circ},180^{\circ}),(90^{\circ},270^{\circ})\}$. When including only frequencies in the sensitivity band of current gravitational-wave detectors, there is no significant difference in the time series of the gravitational-wave signals at different observer angles. However, when frequencies below 10 Hz are included, the figure clearly shows a significant difference in amplitude between the different angles.
To determine if the observer angle has a significant impact on the detectability of the gravitational-wave signals, we calculate their maximum detection distances at the 6 different angles for a variety of different gravitational-wave detectors. We define the maximum detectable distance as the distance required for an optimal signal to noise ratio (SNR) of 8. We include the Advanced LIGO detector at design sensitivity, the Cosmic Explorer detector (Evans et al., 2021), the Einstein Telescope (Hild et al., 2011), and the proposed moon-based gravitational-wave detector, GLOC (Jani & Loeb, 2020). The noise curves for the different detectors are shown in Figure 8. For the Einstein Telescope, we use the D configuration (Hild et al., 2011). For the lunar detector, we use the conservative noise curve from Jani & Loeb (2020). Due to our short duration simulations, we consider a minimum frequency of 2 Hz.
The results are shown in Table 2. We show the frequency at which the gravitational-wave amplitude peaks. This is because the gravitational-wave detectors sensitivity is very frequency-dependent, and the central frequency is also measured by low-latency gravitational-wave unmodelled searches. For models SFHo_nr, SFHx_rr, and SFHx_nr, the gravitational-wave amplitude peaks shortly after shock revival, and the frequency at which the peak emission occurs varies by a few hundred Hz depending on the orientation angle, which is similar to our previous models. However, model SFHo_rr is different, as the gravitational-wave signal is fairly uniform in amplitude for the duration of the entire signal, resulting in a much larger variance in the peak frequency. This is similar to, for example, some models in Bugli et al. (2022). The frequency along the f/g-mode at which the gravitational-wave emission peaks can vary a lot between different CCSN simulation codes, as well as within our own simulations.
@@
For the Einstein Telescope, the maximum detection distances are above a Mpc for the rotating models, and a few hundred kpc for the non-rotating models. The detection distances increase by as much as a few hundred kpc when the neutrino component of the gravitational-wave signal is included. This shows that including the gravitational waves due to the asymmetric neutrino emission is important for building an accurate science case for the detection of CCSNe in Einstein Telescope.
For a gravitational-wave detector on the moon, SFHo_rr can be detected out to $\sim 3$ Mpc, SFHo_nr out to $\sim 600$ kpc, SFHx_rr out to $\sim 2$ Mpc, and SFHx_nr out to $\sim 600$ kpc. Due to the better low-frequency sensitivity on the moon, there is a much larger dependence on source orientation, with the detection distances for SFHx_rr differing by over 1 Mpc for different source orientations, which shows that including multiple source orientations is important for detectability studies for CCSNe in low-frequency gravitational-wave detectors. This is certainly an underestimation of the actual maximum distance that we could detect CCSNe on the moon. This is because our waveforms are too short in duration, and can only go down to frequencies of 2 Hz, when the full frequency band of a moon detector goes down to $\sim 0.1$ Hz. The amplitude at low frequencies would also grow further with a longer simulation time.
@@
In Figure 10, we compare the detection distances for our new models in this paper to our models from our previous studies. We include a non-rotating low mass model he3.5 (Powell & Müller, 2019), two typical non-rotating neutrino-driven explosions models s18 and y20 (Powell & Müller, 2019, 2020), a rapidly rotating model with no magnetic fields model m39 (Powell & Müller, 2020) and two models m39_B12 and m39_B10 both with rapid rotation and strong magnetic fields (Powell et al., 2023). We show the maximum detection distance in a single Advanced LIGO detector. There is a clear relationship between the explosion energy of the source and the gravitational-wave detectability. Our new models have comparable detection distances to previous models without magnetic fields. Typical CCSNe have explosion energies of $10^{51}$ erg, but that energy is not reproduced in the majority of 3D CCSN simulations. As our rotating gravitational-wave signals have energies closer to $10^{51}$ erg, they may be more representative of the actual maximum detection distance for typical CCSNe in Advanced LIGO. A network of gravitational-wave detectors, and more complete waveforms, would further enhance the detection distances for CCSNe.
@@
Table 2: For each of our four simulations we show the model name, the six different source angles used, the maximum detection distances in LIGO for the matter component only, the Einstein Telescope for matter only (m) and matter plus neutrinos (m+n), the moon detector for both matter and neutrinos. For each source orientation, we also show the frequency where the gravitational-wave energy is highest. We define the maximum detectable distance as the distance required for an optimal SNR of 8.
@@ 7 Signal interpretation and universal relations
Figure 11: The spectrograms of our gravitational-wave signals, as shown in Figure 6, with our models fit to the universal relation from Torres-Forné et al. (2019) in white, and Sotani et al. (2021) in pink. Both relations are a reasonably good fit for our models, even when the source is rapidly rotating. Top left is SFHo_rr, top right is SFHo_nr, bottom left is SFHx_rr, and bottom right is SFHx_nr.
Many groups have now formulated universal relations to describe the relationship between the gravitational-wave frequency of the modes and properties like the PNS mass, PNS radius and the shock radius (Torres-Forné et al., 2019; Sotani et al., 2021). These universal relations have been used in gravitational-wave studies for the purpose of gravitational-wave signal interpretation (Powell & Müller, 2022; Bruel et al., 2023). However, the current universal relations do not account for rotation or magnetic fields. They also do not account for the different possible physical mechanisms responsible for the gravitational waves (Mezzacappa et al., 2023), which we have not investigated in this study.
@@
Note, however, that the frequencies computed in the pseudo-Newtonian approximation used in our simulations will be systematically too high. The relations may thus provide a good description of the mode trajectory, perhaps after additional calibration. Our results show that some caution should be used when applying universal relations for the interpretation of the source properties from a real CCSN gravitational-wave detection.
To be able to use the universal relations to determine the source parameters of a CCSN, we need to be able to reconstruct the signal from the gravitational-wave data. Current searches for gravitational waves from CCSNe show that a high SNR of $\sim 20$ is needed before the signal can be detected, and waveform reconstruction methods have needed a SNR of $\sim 20$ for the reconstructed waveform to have a 50% overlap with the original waveform (Szczepańczyk et al., 2021; Raza et al., 2022). Therefore, there also needs to be improvements to CCSNe gravitational-wave searches and waveform reconstruction techniques for us to be able to confidently estimate the parameters of the source when a real CCSN gravitational-wave signal occurs.

### v2

@@ 4 Gravitational Wave Properties
Figure 5: The time series of the GW signal from matter for all models as measured for an observer at the equator ($\theta=\phi=90$). The left column shows the plus polarisation, and the right column shows the cross polarisation. The rapidly rotating models have higher GW amplitudes. Some of the models have started to develop low-frequency tails before the end of the simulation time.
The time series of the GWs from matter as measured by an observer in the equatorial plane ($\theta=\phi=90^{\circ}$), are shown in Figure 5. As the core-bounce time was simulated in 2D, the non-rotating models are showing strong ringdown after prompt convection at the start, and slowly evolve to lower GW amplitudes as the initial axisymmetry is lost in favour of less coherent motions in 3D. This has the largest impact on model SFHo_nr, which has a high amplitude of 10 cm at the start in the plus polarisation. This issue does not have an impact on the rotating models where there is a (physically) preferred axis anyway.
The non-rotating models reach GW amplitudes of $\sim 10$ cm. This is similar to typical amplitudes from our previous neutrino-driven explosion models, for example, model s18 from Powell & Müller (2019), and model y20 from Powell & Müller (2020). This shows that the addition of weak magnetic fields does not have a big impact on GW amplitudes. The rotating models reached GW amplitudes of $\sim 40$ cm. This amplitude is similar to our previous rapidly rotating model m39 (Powell & Müller, 2020), which had no magnetic fields, but significantly smaller than our previous models with strong magnetic fields and rapid rotation, models m39_B10 and m39_B12 from Powell et al. (2023). This is consistent with the much lower explosion energy in the current models compared to $\sim 3\times 10^{51}\,\mathrm{erg}$ in m39_B10 and m39_B12, and suggests that a stronger magnetic field strength is required for powerful GW signals from magnetorotational explosions.
@@
Figure 6: Spectrograms of $h_{+}+h_{\times}$ for all our GW signals. The top left panel shows model SFHo_rr, the top right panel shows model SFHo_nr, the bottom left panel shows model SFHx_rr, and the bottom right panel shows model SFHx_nr. The non-rotating models show some low frequency GW emission from the SASI before shock revival time. All models reach high frequencies of above 1000 Hz before the end of the simulation. Model SFHx_rr reaches higher frequencies due to its smaller PNS radius.
Spectrograms of the GW emission are given in Figure 6. At early times, the non-rotating models have a high-power hotspot at low frequency caused by the ringdown after the transition from 2D to 3D. In the non-rotating models, some low-frequency emission due to the SASI is visible before 0.2 s when shock revival stops the SASI activity. The non-rotating models both have a similar high-frequency f/g-mode emission. They both reach a maximum frequency of $\sim 1500$ Hz by the end of the simulation. The GW frequency is higher than in our previous simulations because of the pseudo-Newtonian gravity used in our simulation. Previous studies have shown that the GW frequency is about $\sim 20$% higher than when full general relativity is used (Müller et al., 2013), due to missing relativistic correction terms in the Brunt-Väisälä frequency (including, but not limited to a time dilation factor from the lapse function $\alpha$, see Equations 16 and C8 in Müller et al., 2013) in the pseudo-Newtonian approximation. The rapidly rotating models have no visible lower frequency modes. Their high-frequency emission has a different shape and maximum frequency for the two different EoS, with a slight S-shape for model SFHx_rr. This is due to the difference in the PNS radius between the two models, as shown previously in Figure 1, where the smaller final PNS radius for model SFHx_rr has resulted in a higher maximum frequency of the f/g-mode mode. The dominant emission band is distinctly broader for the models with rapid rotation, with a width of $\sim 500$ Hz instead of $\sim 250$ Hz, reminiscent of the cases with rapid rotation and strong fields in the 2D study of (Jardine et al., 2022). Such a broadening of the dominant emission band may be an indicator for rapid rotation in the event of a GW detection from a Galactic CCSN.
@@ 5 The Neutrino Memory
Figure 7: The GW amplitude time series from both matter and anisotropic neutrino emission. The left column shows the plus polarisation, and the right column shows the cross polarisation, both calculated at $\theta=\phi=90^{\circ}$ (equatorial observer). All models show a significant increase in amplitude at low frequencies (slow variations) compared to the matter signal due to the neutrino emission anisotropy.
In addition to the low-frequency component to the GW emission from matter, there generally is a high-amplitude low-frequency component to the GW emission from the asymmetric emission of neutrinos (neutrino memory, Epstein, 1978). The neutrino memory will usually dominate over the low-frequency matter signal (Marek et al., 2009b; Müller et al., 2013). This component of the GW signal is often ignored in CCSN simulations that focus on GWs, as it is outside of the frequency band for the current GW detectors. However, understanding the GW signals at frequencies below 10 Hz is important for the science case for the next generation of GW detectors. In this section, we show how the GW emission changes when the signal from the asymmetric emission of neutrinos is also included.
To calculate the GW signal from anisotropic emission of neutrinos, we use Equation (24) from Mueller & Janka (1997), which gives the transverse-traceless perturbations $h_{ij}^{\mathrm{TT}}$ of the metric from neutrino memory,
@@
The combined GW emission from both matter and neutrinos is shown in Figure 7. The low-frequency GW signal from asymmetric neutrino emission has significantly higher amplitude than the matter signal alone. All models show a significant increase in GW amplitude when the neutrino memory is included. Model SFHo_rr reaches amplitudes of 200 cm in the plus polarization, model SFHo_nr reaches 140 cm, model SFHx_rr reaches 100 cm, and model SFHo_nr reaches -150 cm. Due to the short simulation time of these models, we can only predict the low-frequency emission down to about 2 Hz, as the minimum measurable frequency is determined by the simulation duration. The fact that the models quickly reach high amplitudes at low frequencies reflects that the models develop large-scale asymmetry in the explosion ejecta soon after the shock is revived. There is slightly less anisotropy in model SFHx_rr than in the others. The GW signal is still growing significantly in time for the non-rotating models at the end of the simulation time. Longer duration simulations by other groups have shown that the neutrino memory can grow to over a 1000 cm at later times (Vartanyan et al., 2023). This shows the need to carry out long duration simulations to fully capture the both the minimum frequency and maximum amplitude of low-frequency GW emission.
We also show the amplitude spectral densities for all models, with the matter components only, and both the matter and neutrino components in Figure 8. Including the neutrino memory results in a much larger amplitude at low frequencies, in the frequency detection bands of Einstein Telescope and proposed moon-based GW detectors.
Figure 8: The amplitude spectral density (ASD) of all models, as measured at a distance of 100 kpc, and the sensitivity curves of the LIGO, Einstein Telescope (ET), Cosmic Explorer and moon-based (GLOC) GW detectors. The left panel shows the matter component only. The right panel shows the spectrum from the combined matter and neutrino signal. The GW spectrum from matter peaks at frequencies of above 1000 Hz, and the GW spectrum from asymmetric neutrino emission peaks below 10 Hz.
@@ 6 Gravitational Wave Directional Dependence and detectability
Figure 9: The GW time series of all four models as measured at six different observer angles. (top left) The GWs for model SFHo_rr. (top right) The GWs for model SFHo_nr. (bottom left) The GWs for model SFHx_rr. (bottom right) The GWs for model SFHx_nr. The amplitude of the low-frequency GW emission can vary significantly at different source angles.
In Figure 9, we show the GW emission calculated at six different representative source angles $(\theta,\phi)\in\{(0^{\circ},0^{\circ}),(180^{\circ},0^{\circ}),(90^{\circ},0^{\circ}),(90^{\circ},90^{\circ}),(90^{\circ},180^{\circ}),(90^{\circ},270^{\circ})\}$. When including only frequencies in the sensitivity band of current GW detectors, there is no significant difference in the time series of the GW signals at different observer angles. However, when frequencies below 10 Hz are included, the figure clearly shows a significant difference in amplitude between the different angles.
To determine if the observer angle has a significant impact on the detectability of the GW signals, we calculate their maximum detection distances at the 6 different angles for a variety of different GW detectors. We define the maximum detectable distance as the distance required for an optimal signal to noise ratio (SNR) of 8. We include the Advanced LIGO detector at design sensitivity, the Cosmic Explorer detector (Evans et al., 2021), the Einstein Telescope (Hild et al., 2011), and the proposed moon-based GW detector, GLOC (Jani & Loeb, 2020). The noise curves for the different detectors are shown in Figure 8. For the Einstein Telescope, we use the D configuration (Hild et al., 2011). For the lunar detector, we use the conservative noise curve from Jani & Loeb (2020). Due to our short duration simulations, we consider a minimum frequency of 2 Hz.
The results are shown in Table 2. We show the frequency at which the GW amplitude peaks. This is because the GW detectors sensitivity is frequency-dependent, and the central frequency is also measured by low-latency GW unmodelled searches. For models SFHo_nr, SFHx_rr, and SFHx_nr, the GW amplitude peaks shortly after shock revival, and the frequency at which the peak emission occurs varies by a few hundred Hz depending on the orientation angle, which is similar to our previous models. However, model SFHo_rr is different, as the GW signal is fairly uniform in amplitude for the duration of the entire signal, resulting in a much larger variance in the peak frequency. This is similar to, for example, some models in Bugli et al. (2022). The frequency along the f/g-mode at which the GW emission peaks can vary a lot between different CCSN simulation codes, as well as within our own simulations.
@@
For the Einstein Telescope, the maximum detection distances are above a Mpc for the rotating models, and a few hundred kpc for the non-rotating models. The detection distances increase by as much as a few hundred kpc when the neutrino component of the GW signal is included. This shows that including the GWs due to the asymmetric neutrino emission is important for building an accurate science case for the detection of CCSNe in Einstein Telescope.
For a GW detector on the moon, SFHo_rr can be detected out to $\sim 3$ Mpc, SFHo_nr out to $\sim 600$ kpc, SFHx_rr out to $\sim 2$ Mpc, and SFHx_nr out to $\sim 600$ kpc. Due to the better low-frequency sensitivity on the moon, there is a much larger dependence on source orientation, with the detection distances for SFHx_rr differing by over 1 Mpc for different source orientations, which shows that including multiple source orientations is important for detectability studies for CCSNe in low-frequency GW detectors. This is certainly an underestimation of the actual maximum distance that we could detect CCSNe on the moon. This is because our waveforms are too short in duration, and can only go down to frequencies of 2 Hz, when the full frequency band of a moon detector goes down to $\sim 0.1$ Hz. The amplitude at low frequencies would also grow further with a longer simulation time.
@@
In Figure 10, we compare the detection distances for our new models in this paper to our models from our previous studies. We include a non-rotating low mass model he3.5 (Powell & Müller, 2019), two typical non-rotating neutrino-driven explosions models s18 and y20 (Powell & Müller, 2019, 2020), a rapidly rotating model with no magnetic fields model m39 (Powell & Müller, 2020) and two models m39_B12 and m39_B10 both with rapid rotation and strong magnetic fields (Powell et al., 2023). We show the maximum detection distance in a single Advanced LIGO detector. There is a clear relationship between the explosion energy of the source and the GW detectability. Our new models have comparable detection distances to previous models without magnetic fields. Typical CCSNe have explosion energies of $10^{51}$ erg, but that energy is not reproduced in the majority of 3D CCSN simulations. As our rotating GW signals have energies closer to $10^{51}$ erg, they may be more representative of the actual maximum detection distance for typical CCSNe in Advanced LIGO. A network of GW detectors, and more complete waveforms, would further enhance the detection distances for CCSNe.
@@
Table 2: For each of our four simulations we show the model name, the six different source angles used, the maximum detection distances in LIGO for the matter component only, the Einstein Telescope for matter only (m) and matter plus neutrinos (m+n), the moon detector for both matter and neutrinos. For each source orientation, we also show the frequency where the GW energy is highest. We define the maximum detectable distance as the distance required for an optimal SNR of 8.
@@ 7 Signal interpretation and universal relations
Figure 11: The spectrograms of our GW signals, as shown in Figure 6, with our models fit to the universal relation from Torres-Forné et al. (2019) in white, and Sotani et al. (2021) in pink. Both relations are a reasonably good fit for our models, even when the source is rapidly rotating. Top left is SFHo_rr, top right is SFHo_nr, bottom left is SFHx_rr, and bottom right is SFHx_nr.
Many groups have now formulated universal relations, that are designed to formulate the relationship between the GW frequency of the modes and properties like the PNS mass, PNS radius and the shock radius for every CCSN GW signal (Torres-Forné et al., 2019; Sotani et al., 2021). These universal relations have been used in GW studies for the purpose of GW signal interpretation (Powell & Müller, 2022; Bruel et al., 2023). However, the current universal relations do not account for rotation or magnetic fields. They also do not account for the different possible physical mechanisms responsible for the GWs (Mezzacappa et al., 2023), which we have not investigated in this study.
@@
Note, however, that the frequencies computed in the pseudo-Newtonian approximation used in our simulations will be systematically too high. The relations may thus provide a good description of the mode trajectory, perhaps after additional calibration. Our results show that some caution should be used when applying universal relations for the interpretation of the source properties from a real CCSN GW detection.
To be able to use the universal relations to determine the source parameters of a CCSN, we need to be able to reconstruct the signal from the GW data. Current searches for GWs from CCSNe show that a high SNR of $\sim 20$ is needed before the signal can be detected, and waveform reconstruction methods have needed a SNR of $\sim 20$ for the reconstructed waveform to have a 50% overlap with the original waveform (Szczepańczyk et al., 2021; Raza et al., 2022). Therefore, there also needs to be improvements to CCSNe GW searches and waveform reconstruction techniques for us to be able to confidently estimate the parameters of the source when a real CCSN GW signal occurs.